# Empirical Bayes Estimation in Heterogeneous Coefficient Panel Models

### v1

@@ Abstract
We develop an empirical Bayes (EB) G-modeling framework for short-panel linear models with multidimensional heterogeneity and nonparametric prior. Specifically, we allow heterogeneous intercepts, slopes, dynamics, and a non-spherical error covariance structure. We establish identification and consistency of the nonparametric maximum likelihood estimator (NPMLE) under general conditions, and provide low-level sufficient conditions for several models of empirical interest. Conditions for regret consistency of the resulting EB estimators are also established. The NPMLE is computed using a Wasserstein-Fisher-Rao gradient flow algorithm adapted to panel regressions. Using data from the Panel Study of Income Dynamics, we find that the slope coefficient for potential experience is substantially heterogeneous and negatively correlated with the random intercept, and that error variances and autoregressive coefficients vary significantly across individuals. The EB estimates reduce mean squared prediction errors relative to individual maximum likelihood estimates.
@@ 1 Introduction
where $a_{i}\in\mathbb{R}$ is a random intercept, $b_{i}\in\mathbb{R}$ a random slope, $\rho_{i}\in(-1,1)$ a random autoregressive coefficient, $\sigma_{i}^{2}>0$ a random variance, and $u_{i0}$ follows the stationary distribution $\mathcal{N}(0,\sigma_{i}^{2}/(1-\rho_{i}^{2}))$ conditional on $(\sigma_{i}^{2},\rho_{i})$. We refer to the model in (1) as the HIVDX model since it admits heterogeneity in four dimensions: intercepts ($I$), error variances ($V$), dynamics ($D$) through the autoregressive coefficient, and the coefficient on the covariate $X_{2}$. The widely studied Gaussian location-scale model $Y_{it}=a_{i}+\sigma_{i}e_{it}$, as well as models that impose $\rho_{i}=\rho$ and $b_{i}=b$, are special cases of HIVDX.
@@
We analyze the HIVDX model and the special cases in Section 2 using a general framework that partitions the random parameters into two sets: regression coefficients $\beta_{i}=(a_{i},b_{i})$ of dimension $d_{\beta}=2$, and covariance parameters $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$ of dimension $d_{\delta}=2$. The parameters can be estimated separately for each $i$ by maximum likelihood. Instead, we assume that the random parameters $\theta_{i}\equiv(\beta_{i},\delta_{i})$ are drawn from a true but unknown common distribution $G_{*}$. Viewing $G_{*}$ as a prior, the optimal decision rule for each unit under compound squared-error loss is the posterior mean ${\theta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\theta_{i}\mid Y_{i},X_{2,i}]$, where $Y_{i}=(Y_{i1},\ldots,Y_{iT})^{\prime}$ is a $T\times 1$ vector of outcomes, $X_{2,i}=(X_{2,i1},\ldots,X_{2,iT})^{\prime}$ is a vector of observed covariates, and $\mathbb{E}_{G_{*}}[\cdot]$ denotes expectation with respect to $\theta_{i}\sim G_{*}$. We refer to $\theta_{i}^{*}$ as the oracle estimator.
@@
Specifying that $\theta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}G_{*}$ is grounded in the principle of empirical Bayes (EB) introduced by Robbins (1956). This approach is intrinsically frequentist because it estimates the prior distribution from the data rather than specifying it. EB estimation improves on individual-level maximum likelihood estimation through shrinkage, borrowing strength across units to learn both the individual parameters $\theta_{i}$ and their population distribution $G_{*}$. Following the terminology of Efron (2014) the procedure is referred to as G-modeling when $G_{*}$ is estimated directly. The resulting EB estimator is given by $\hat{\theta}^{\mathrm{EB}}_{i}:=\mathbb{E}_{\hat{G}}[\theta_{i}\,|\,Y_{i},X_{2,i}]$, where $\hat{G}$ denotes an estimator of $G_{*}$. An EB estimator $\hat{\tau}_{i}^{{\mathrm{EB}}}(\theta_{i})=\mathbb{E}_{\hat{G}}[\tau_{i}\,|\,Y_{i},X_{2,i}]$ for general unit-specific parameters $\tau_{i}$ defined as functions of $\theta_{i}$ can also be defined. While parametric G-modeling is convenient, nonparametric modeling guards against misspecification. Recent reviews of EB and its applications in economics include Koenker and Gu (2024), Koenker and Gu (2025), and Walters (2024). To our knowledge, no nonparametric EB analysis has considered all four sources of heterogeneity present in the HIVDX model.
@@
We study identification, estimation and and optimality of the HIVDX model within an EB framework. The generality of our model with $d_{\beta}>1$ and $d_{\delta}>0$ introduces several challenges. First, while assuming the identification of $G_{*}$ in the Gaussian location model with i.i.d. errors without formal proof may be defensible, this assumption cannot be taken for granted in the more complex settings. Theorem 1 establishes generally that identification requires: (i) the design matrix to have full column rank within its support, and (ii) the covariance structure to remain identifiable after removing the mean component by appropriately differencing the data. These conditions imply that $d_{\beta}<T$, leaving $T-d_{\beta}$ degrees of freedom to identify the covariance structure, thereby imposing a constraint on the number of heterogeneous variance parameters ($d_{\delta}$). Proposition 1 applies Theorem 1 to the location model ($d_{\beta}=1$) and illustrates how differencing facilitates identification. Specifically, we show that $T\geq 3$ suffices for AR(1) errors, whereas ARMA(1,1) errors require $T\geq 4$. The analysis is, however, more involved when $d_{\beta}>1$, as the mean and variance components are convolved within the outcome distribution. Now whether and how $G_{*}$ can be identified depends critically on the specific time-series variation of the covariate. Proposition 3 establishes that if $X_{2,it}$ is such that $\Delta X_{2,i2}=\pm\Delta X_{2,i3}\neq 0$ or $\Delta X_{2,i2}\neq\Delta X_{2,i4}$, where $\Delta$ denotes the first-difference operator, then $G_{*}$ is nonparametrically identified with $T=4$. However, Proposition 4 shows that identification is possible with $T\geq 5$ for a different $X_{2,it}$ satisfying the weaker condition that $\Delta X_{2,it}\neq 0$ for some $t=2,\ldots,T$.
@@
The second challenge lies in establishing the consistency of the HIVDX model. The classic proof of Kiefer and Wolfowitz (1956) assumes compactness, but when $\beta_{i}$ is allowed to be unbounded, the space of mixing distributions is non-compact. While the case of ${d_{\beta}}=1$ can be resolved by embedding $\mathbb{R}$ into its compact completion $\infty$$\infty$], the multidimensional setting (${d_{\beta}}\geq 2$) presents unique difficulties because mass can diverge along infinitely many distinct directions. To address this, Section 4 adopts a metric induced by the vague topology to effectively compactify the space of prior distributions. Theorem 2 establishes the almost sure consistency of the NPMLE with respect to this metric. Building on this result, Theorem 3 in Section 5 provides conditions for the regret consistency (equivalently, asymptotic optimality) of the EB estimator for $\tau_{i}$, taking into account that $\tau_{i}$ may involve functions of the random coefficients and the data. For example, the oracle one-step-ahead prediction under quadratic loss is ${\tau}^{*}_{i}=\mathbb{E}_{G_{*}}[a_{i}+b_{i}X_{2,iT+1}+\rho_{i}Y_{iT}\,|\,Y_{i},X_{i},X_{2,iT+1}]-\mathbb{E}_{G_{*}}[\rho_{i}a_{i}+\rho_{i}b_{i}X_{2,iT}\,|\,Y_{i},X_{i}]$. The prediction error $\hat{\tau}^{{\mathrm{EB}}}_{i}-\tau_{i}^{*}$ thus involves interaction of the errors in estimating the random coefficients and $(Y_{i},X_{i})$. Proposition 5 provides low-level sufficient conditions for the EB prediction to be regret-consistent.
The third challenge arises from the computational complexity associated with multiple sources of heterogeneity. Building on the interior-point method of Koenker and Mizera (2014), Gu and Koenker (2017) discretize $G_{*}$ over a fixed grid and solve the dual formulation of the NPMLE via convex optimization. While effective in one or two dimensions, these methods do not scale well as the dimension of the parameter space increases. We also approximate $G_{*}$ using a discrete distribution, but we employ an algorithm capable of flexibly handling higher-dimensional heterogeneity. To this end, Section 6 adapts the Wasserstein-Fisher-Rao (WFR) gradient flow algorithm of Yan, Wang, and Rigollet (2024) to the panel regression setting. In this algorithm, the Wasserstein step adaptively relocates support points to accelerate convergence, while the Fisher-Rao step updates the probability weights via Bayes’ rule.
@@
We use the EB approach to analyze the income data in the Panel Study of Income Dynamics (PSID) in Section 7. The model allows earnings $Y_{it}$ to have unit-specific mean $a_{i}$ and response $b_{i}$ to experience $X_{2,it}$, with a heteroskedastic error covariance structure that depends on internal dynamics $\rho_{i}$ and exposure to shocks through $\sigma_{i}^{2}$. We find that (i) the slope coefficient $b_{i}$ for potential experience exhibits substantial heterogeneity and is negatively correlated with the random intercept $a_{i}$, and (ii) there is pronounced cross-sectional heterogeneity in both the variance ($\sigma_{i}^{2}$) and the dynamics ($\rho_{i}$). The EB estimates reduce mean squared prediction errors relative to individual maximum likelihood estimates. Section 8 presents results from Monte Carlo experiments calibrated to the application.
@@
Before turning to the main part of the paper, we briefly discuss recent work that, while not directly related to our approach, are conceptually connected to the broader EB literature. Chen (2024) analyzes conditionally Gaussian settings with known variances $\sigma_{i}^{2}$, constructing studentized data by estimating the conditional mean and variance of outcomes given $\sigma_{i}^{2}$, and then estimating the prior for use in EB decision rules. The paper emphasizes the importance of nonparametric estimation in both stages and illustrates the method using data on economic mobility. In contrast, we allow for unobserved $\sigma_{i}^{2}$ and heterogeneous dynamics across multiple time periods.
Kwon (2025) considers a multivariate model in which $Y_{i}$ given $\theta_{i}\in\mathbb{R}^{T}$ follows $\mathcal{N}(\theta_{i},\Sigma_{i})$ with known covariance matrices $\Sigma_{i}$. Under the hierarchical specification $\theta_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\Lambda)$, the shrinkage estimator $\Lambda(\Lambda+\Sigma_{i})^{-1}Y_{i}$ is shown to be optimal when $\Lambda$ minimizes Stein’s unbiased risk estimate, without requiring normality for validity. Our framework differs in that $\Sigma_{i}$ is unobserved and we allow for random slopes on regressors, whereas Kwon (2025) assumes homogeneous coefficients and removes the regression component prior to estimation. Cheng, Ho, and Schorfheide (2025) propose a parametric EB estimator for two-way effects in linked data, using a prior that captures patterns of assortative matching. Their framework is more related to Kwon (2025) than ours in that the shrinkage parameters are chosen by minimizing an unbiased risk estimate. Both studies are applied to teacher value-added models.
@@
Gaillac (2024) develops posterior mean estimators for cross-sectional random coefficient models in which the coefficients are independent of the regressors. In a panel extension, the paper constructs time averages for each unit and allows heterogeneity only in time-invariant regressors, while assuming homogeneous coefficients for time-varying regressors. In contrast, our framework focuses on random coefficients associated with regressors that vary over time.
@@ 2 The Econometric Setup and EB Modeling
For the estimation of $(\theta_{i})_{i=1}^{N}$, two conceptual frameworks are available. The separate decision framework treats each $\theta_{i}$ as fixed and minimizes the individual risk $\mathbb{E}[\|\hat{\theta}_{i}-\theta_{i}\|^{2}\,|\,\theta_{i}]$. This yields the individual MLE,
@@
which provides an asymptotically efficient estimator for $\theta_{i}$ as $T\to\infty$. However, Stein (1956) and James and Stein (1961) demonstrate that the MLE is inadmissible under squared-error loss when the dimension of the parameter vector is at least three (here, $N\geq 3$). In this setting, the MLE is dominated by a shrinkage estimator that achieves lower total risk by pooling information across observations.
The second approach, the compound decision framework, treats $(\theta_{i})_{i=1}^{N}$ as i.i.d. draws from a common distribution $G_{*}$ and seeks to minimize the compound risk $\mathbb{E}_{\theta_{i}\sim G_{*}}[\|\hat{\theta}_{i}-\theta_{i}\|^{2}]$. Viewing $G_{*}$ as a prior, the optimal decision rule for each unit is the posterior mean, defined in the introduction as the oracle estimator:
${\theta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\theta_{i}\,|\,Y_{i},X_{i}]=\frac{\int_{\Theta}\theta\ell(Y_{i}\,|\,X_{i},\theta)dG_{*}(\theta)}{f_{G_{*}}(Y_{i},X_{i})}.$ (4)
@@
F-modeling typically relies on the assumption that the likelihood belongs to the linear exponential family, offering computational convenience when this condition holds. Walters (2024) provides a survey of F-modeling applications within labor economics. Liu, Moon, and Schorfheide (2020) estimate the posterior mean of $a_{i}$ by F-modeling using Tweedie’s formula. However, this method does not readily extend to the HIVDX model. We adopt the alternative of G-modeling, which explicitly estimates the unknown prior $G_{*}$. Shen and Wu (2025) demonstrate that G-modeling has superior statistical properties in Poisson mixture models than F-modeling.
@@
where the empirical criterion is given by $F_{N}(G):=N^{-1}\sum_{i=1}^{N}\log f_{G}(Y_{i},X_{i})$. We refer to $\hat{G}$ as the nonparametric maximum likelihood estimator (NPMLE) of $G_{*}$. Substituting $\hat{G}$ for $G_{*}$ in (4) yields the G-modeling EB estimator for $\theta_{i}$:

### v2

@@ Abstract
We develop an empirical Bayes (EB) G-modeling framework for short-panel linear models with nonparametric prior for the random intercepts, slopes, dynamics, and non-spherical error variances. We establish identification and consistency of the nonparametric maximum likelihood estimator (NPMLE) under general conditions, and provide low-level sufficient conditions for several models of empirical interest. Conditions for regret consistency of the EB estimators are also established. The NPMLE is computed using a Wasserstein-Fisher-Rao gradient flow algorithm adapted to panel regressions. Using data from the Panel Study of Income Dynamics, we find that the slope coefficient for potential experience is substantially heterogeneous and negatively correlated with the random intercept, and that error variances and autoregressive coefficients vary significantly across individuals. The EB estimates reduce mean squared prediction errors relative to individual maximum likelihood estimates.
@@ 1 Introduction
where $a_{i}\in\mathbb{R}$ is a random intercept, $b_{i}\in\mathbb{R}$ a random slope, $\rho_{i}\in(-1,1)$ a random autoregressive coefficient, $\sigma_{i}^{2}>0$ a random variance, and $u_{i0}$ follows the stationary distribution $\mathcal{N}(0,\sigma_{i}^{2}/(1-\rho_{i}^{2}))$ conditional on $(\sigma_{i}^{2},\rho_{i})$. We refer to the model in (1) as HIVDX since it admits heterogeneity in four dimensions: intercepts ($I$), error variances ($V$), dynamics ($D$) through the autoregressive coefficient, and the coefficient on the exogenous covariate $X_{2}$. The widely studied Gaussian location-scale model $Y_{it}=a_{i}+\sigma_{i}e_{it}$, as well as models that impose $\rho_{i}=\rho$ and $b_{i}=b$, are special cases of HIVDX.
@@
We analyze the HIVDX model and the special cases in Section 2 using a general framework that partitions the random parameters into two sets: regression coefficients $\beta_{i}=(a_{i},b_{i})$ of dimension $d_{\beta}=2$, and covariance parameters $\delta_{i}=(\sigma_{i}^{2},\rho_{i})$ of dimension $d_{\delta}=2$. The parameters can be estimated separately for each $i$ by maximum likelihood. Instead, we assume that the random parameters $\theta_{i}\equiv(\beta_{i},\delta_{i})$ are drawn from a true but unknown common distribution $G_{*}$. Viewing $G_{*}$ as a prior, the optimal decision rule for each unit under compound squared-error loss is the posterior mean ${\theta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\theta_{i}\mid Y_{i},X_{2,i}]$, where $Y_{i}=(Y_{i1},\ldots,Y_{iT})^{\prime}$, $X_{2,i}=(X_{2,i1},\ldots,X_{2,iT})^{\prime}$, and $\mathbb{E}_{G_{*}}[\cdot]$ denotes expectation with respect to $\theta_{i}\sim G_{*}$.
@@
Specifying that $\theta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}G_{*}$ is grounded in the principle of empirical Bayes (EB) introduced by Robbins (1956). This approach is intrinsically frequentist because it estimates the prior distribution from the data. EB estimation improves on individual-level maximum likelihood estimation through shrinkage, borrowing strength across units to learn both the individual parameters $\theta_{i}$ and their population distribution $G_{*}$. Following the terminology of Efron (2014), the procedure is referred to as G-modeling when $G_{*}$ is estimated directly. The resulting EB estimator is given by $\hat{\theta}^{\mathrm{EB}}_{i}:=\mathbb{E}_{\hat{G}}[\theta_{i}\,|\,Y_{i},X_{2,i}]$, where $\hat{G}$ denotes an estimator of $G_{*}$. An EB estimator $\hat{\tau}_{i}^{{\mathrm{EB}}}(\theta_{i})=\mathbb{E}_{\hat{G}}[\tau_{i}\,|\,Y_{i},X_{2,i}]$ for general unit-specific parameters $\tau_{i}$ defined as functions of $\theta_{i}$ can also be defined. While parametric G-modeling is convenient, nonparametric modeling guards against misspecification. Recent reviews of EB and its applications in economics include Koenker and Gu (2024), Koenker and Gu (2025), and Walters (2024). To our knowledge, no nonparametric EB analysis has considered all four sources of heterogeneity present in the HIVDX model.
@@
We study identification and estimation of the HIVDX model within an EB framework. The generality of our model with $d_{\beta}>1$ and $d_{\delta}>0$ introduces several challenges. First, while assuming the identification of $G_{*}$ in the Gaussian location model with i.i.d. errors without formal proof may be defensible, this assumption cannot be taken for granted in more complex settings. Theorem 1 establishes that identification requires: (i) the design matrix to have full column rank almost surely, and (ii) the covariance structure to remain identifiable after removing the mean component by appropriately differencing the data. These conditions imply that $d_{\beta}<T$, leaving $T-d_{\beta}$ degrees of freedom to identify the covariance structure; this imposes a constraint on the dimension of heterogeneous variance parameters ($d_{\delta}$).
We apply Theorem 1 to the location model ($d_{\beta}=1$) and illustrate how differencing facilitates identification. Specifically, we show that $T\geq 3$ suffices for AR(1) errors (Proposition 1), whereas ARMA(1,1) errors require $T\geq 4$ (Proposition 2). The analysis is, however, more involved when $d_{\beta}>1$, as the mean and variance components are convolved within the outcome distribution. Now whether and how $G_{*}$ can be identified depends critically on the specific time-series variation of the covariate. We examine two canonical examples. The first example includes models with a common time trend and heterogeneous initial conditions (e.g., potential experience in income dynamics) as a special case. For this example, Proposition 3 establishes that $G_{*}$ is nonparametrically identified with $T\geq 4$. The second example includes a common mean shift as a special case. For this case, Proposition 4 shows that identification of $G_{*}$ is possible with $T\geq 5$. Identification must be worked out on a case by case basis.
@@
The second challenge lies in establishing consistent estimation of the HIVDX model. The classic proof of Kiefer and Wolfowitz (1956) assumes compactness, but when $\beta_{i}$ is allowed to be unbounded, the space of mixing distributions is non-compact. While the case of ${d_{\beta}}=1$ can be resolved by embedding $\mathbb{R}$ into its compact completion $\infty$$\infty$], the multidimensional setting (${d_{\beta}}\geq 2$) presents unique difficulties because mass can diverge along infinitely many distinct directions. To address this, Section 4 adopts a metric induced by the vague topology to effectively compactify the space of prior distributions. Theorem 2 establishes almost sure consistency of the nonparametric maximum likelihood estimator (NPMLE) with respect to this metric. Building on this result, Theorem 3 in Section 5 provides conditions for regret consistency (equivalently, asymptotic optimality) of the EB estimator for $\tau_{i}$, taking into account that $\tau_{i}$ may involve functions of the random coefficients and the data. For example, the optimal (but infeasible) one-step-ahead prediction under quadratic loss is ${\tau}^{*}_{i}=\mathbb{E}_{G_{*}}[a_{i}+b_{i}X_{2,iT+1}+\rho_{i}Y_{iT}\,|\,Y_{i},X_{i},X_{2,iT+1}]-\mathbb{E}_{G_{*}}[\rho_{i}a_{i}+\rho_{i}b_{i}X_{2,iT}\,|\,Y_{i},X_{i}]$. The prediction error $\hat{\tau}^{{\mathrm{EB}}}_{i}-\tau_{i}^{*}$ thus involves interaction of the errors in estimating the random coefficients and $(Y_{i},X_{i})$. Proposition 5 provides low-level sufficient conditions for the EB prediction to be regret-consistent.
The third challenge arises from the computational complexity associated with multiple sources of heterogeneity. In location models, the NPMLE of $G$ is typically computed by optimizing over a fixed grid of $m$ points. The location of the grid points and the size of the grid are usually fixed during the iterations. However, as $d_{\theta}$ increases, the grid required to maintain accuracy needs to grow exponentially. It would be desirable to have an algorithm that allows the support points to move adaptively. This motivates a shift from Euclidean optimization to optimization on the space of probability measures, where the updates are governed by gradient flows. In this framework, the choice of geometry is critical. For example, the Fisher-Rao gradient flow corresponds to reweighting fixed support points, whereas the Wasserstein gradient flow transports mass across the parameter space, effectively moving the support. Recently, Yan, Wang, and Rigollet (2024) demonstrate the theoretical and computational advantages of combining these geometries for the Gaussian location mixture model. Motivated by this encouraging result, Section 6 develops a Wasserstein-Fisher-Rao (WFR) gradient flow algorithm for panel regressions with multiple sources of heterogeneity.
@@
We use the EB approach to analyze the income data in the Panel Study of Income Dynamics (PSID) in Section 7. The model allows earnings $Y_{it}$ to have unit-specific mean $a_{i}$ and response $b_{i}$ to potential experience $X_{2,it}$, with a heteroskedastic error covariance structure that depends on internal dynamics $\rho_{i}$ and exposure to shocks through $\sigma_{i}^{2}$. We find that (i) the slope coefficient $b_{i}$ for potential experience exhibits substantial heterogeneity and is negatively correlated with the random intercept $a_{i}$, and (ii) there is pronounced cross-sectional heterogeneity in both the variance ($\sigma_{i}^{2}$) and the dynamics ($\rho_{i}$). The EB estimates reduce mean squared prediction errors relative to individual maximum likelihood estimates. Section 8 shows that features of the EB estimates and forecast error reduction are replicated in Monte Carlo experiments calibrated to the application.
@@
Before turning to the main part of the paper, we briefly discuss recent work that, while not directly related to our approach, are conceptually connected to the broader EB literature. Chen (2024) analyzes conditionally Gaussian settings with known variances $\sigma_{i}^{2}$. He first studentized the data by estimating the conditional mean and variance of outcomes given $\sigma_{i}^{2}$, and then estimate the prior to plug into EB decision rules. The paper emphasizes the importance of nonparametric estimation in both stages. In contrast, we allow for unobserved $\sigma_{i}^{2}$ and heterogeneous dynamics across multiple time periods.
@@
Kwon (2025) considers a multivariate Gaussian model with a hierarchical prior, while Cheng, Ho, and Schorfheide (2025) propose a parametric EB estimator for two-way effects with assortative matching. Both methods rely on linear shrinkage rules selected to minimize unbiased risk estimates. In contrast, we adopt a fully nonparametric EB approach via NPMLE. Gaillac (2024) develops posterior mean estimators for multiple random coefficient models for cross-sectional regressions. In a panel extension, the paper constructs time averages for each unit and allows heterogeneity only in time-invariant regressors, while assuming homogeneous coefficients for time-varying regressors. In contrast, our framework focuses on random coefficients associated with regressors that vary over time.
@@ 2 The Econometric Setup and EB Modeling
For the estimation of $(\theta_{i})_{i=1}^{N}$, two conceptual frameworks are available. The first is the separate decision framework that treats each $\theta_{i}$ as fixed and minimizes the individual risk $\mathbb{E}[\|\hat{\theta}_{i}-\theta_{i}\|^{2}\,|\,\theta_{i}]$. This yields the individual MLE,
@@
which provides an asymptotically efficient estimator for $\theta_{i}$ as $T\to\infty$. However, Stein (1956) and James and Stein (1961) demonstrate for independent Gaussian observations that the MLE is inadmissible under squared-error loss when the dimension of the parameter vector is at least three (here, $N\geq 3$). In this setting, the MLE is dominated by a shrinkage estimator that achieves lower total risk by pooling information across observations.
The second approach is the compound decision framework. Unlike the separate framework, which minimizes risk for each unit individually, this approach seeks to minimize the aggregate risk over the ensemble of $N$ units. While the strict compound decision formulation treats $(\theta_{i})_{i=1}^{N}$ as fixed, Robbins (1956) demonstrates that this problem can be effectively solved by adopting an EB perspective. By modeling the parameters as i.i.d. draws from a common latent distribution $G_{*}$, we can derive an optimal (oracle) decision rule, which takes the form of the posterior mean:
${\theta}^{*}_{i}:=\mathbb{E}_{G_{*}}[\theta_{i}\,|\,Y_{i},X_{i}]=\frac{\int_{\Theta}\theta\ell(Y_{i}\,|\,X_{i},\theta)dG_{*}(\theta)}{\int_{\Theta}\ell(Y_{i}\,|\,X_{i},\theta)dG_{*}(\theta)}.$ (4)
@@
F-modeling typically relies on the assumption that the likelihood belongs to the linear exponential family, offering computational convenience when this condition holds. Walters (2024) provides a survey of F-modeling applications in labor economics. Liu, Moon, and Schorfheide (2020) estimate the posterior mean of $a_{i}$ by F-modeling using Tweedie’s formula. However, this method does not readily extend to the HIVDX model. We adopt the alternative of G-modeling and explicitly estimate the unknown prior $G_{*}$. Shen and Wu (2025) demonstrate that G-modeling has superior statistical properties in Poisson mixture models than F-modeling.
@@
where the empirical criterion is given by $F_{N}(G):=N^{-1}\sum_{i=1}^{N}\log f_{G}(Y_{i},X_{i})$. We refer to $\hat{G}$ as the NPMLE of $G_{*}$. Substituting $\hat{G}$ for $G_{*}$ in (4) yields the G-modeling EB estimator for $\theta_{i}$: